One apt. is directly above second apt. Downstairs resident calls his upstairs neighbor and says;"If one of you is willing to come downstairs, we'll have the same number of people in both apts." His neighbor responded with; "Why don't one of you come up? Then we,ll have twice as many people up here as you have down there." How many people in each apt? I know the answers, but am having trouble setting it up into a system of linear equations. Can somebody please help me? Thank you so much in advance.
Answer by scott8148(6628)

If one of you is willing to come downstairs, we'll have the same number of people in both apts
... D+1=U-1

Why don't one of you come up? Then we'll have twice as many people up here as you have down there
... 2(D-1)=U+1